### v3

Abstract
Large Language Models (LLMs) are increasingly deployed as interacting agents, forming “LLM societies”. Understanding whether these societies can self-organize and coordinate on norms without external influence is crucial to understand their risks and opportunities. Here we explore their opinion dynamics finding that it is governed by a majority force coefficient such that LLM societies can spontaneously reach consensus only up to a critical group size. This critical size grows exponentially with the language understanding capabilities of the models, exceeding the typical size of informal human groups for advanced LLMs. These results reveal emerging self-organization properties in LLM societies and provide insights for designing collaborative AI systems where coordination is either a goal or a risk.
Introduction
Large Language Models (LLMs) have proven individual capabilities for a wide range of applications, such as summarization [1], sentiment analysis [2, 3], scientific research [4] or mathematical reasoning [5]. Agents driven by LLMs can be used in group settings where several agents interact with each other in collaborative tasks [6, 7, 8]. Collaboration setups where multiple LLMs have different roles and tasks, such as AutoGPT  [^1] https://github.com/Significant-Gravitas/AutoGPT and more recently Microsoft’s AutoGen [9] or OpenAI SWARM [^2] https://github.com/openai/swarm, are qualitatively different from ensembling techniques, where interaction between different models is absent [10]. Recent advancement on on-device LLMs are also leading to AI-powered devices and assistants, such as Siri [^3] https://openai.com/index/openai-and-apple-announce-partnership/ or the Humane AI Pin, that can perform everyday tasks in interaction with each other, for example coordinating events or negotiating prices.
As we move towards a society where LLM agents interact with each other in our behalf and can coordinate, it becomes important to understand their ability to agree with each other in large groups. This can motivate new applications but also help identifying risks stemming from undesired collective behavior. For example, trading bots interacting through the stock market can lead to flash crashes [11]. Current research on the behavior of LLMs has mostly focused on their behavior in isolation [12, 13, 14, 15, 16, 17] and collective behavior has been explored less [18, 19, 20], mostly with a focus on social simulation of network structures [21, 22, 23], opinion and information spreading [24, 25] and online interaction [26, 27]. To be prepared for large numbers of interacting LLM agents, we need to understand if they can display emerging consensus, what determines the abilities of LLM agents to coordinate, and at what scale this can happen.
Group coherence is related to coordination, as the agreement on shared norms creates a social contract that regulates behavior [28]. Norms are also a necessity even in situations where there is no information about the quality or utility of any option, for example in the case of animal coordination when moving collectively [29]. Across species, this leads to a scaling of the average group size with brain structure, with human groups reaching sizes between 150 and 300 [30, 31], as documented by archaeological records [32] and contemporary experience [33]. To reach larger scales, human societies have built institutions and other ways of decision making, but the cognitive limit of about 250 contacts remains even in an online society [34, 35]. In this paradigm, language can work as a tool to learn about the behavior of others in a more efficient way than through observation, for example through gossiping [36, 37], thus leading to the hypothesis that language abilities are a factor in the capability of LLM societies to reach consensus. These are the new questions of AI anthropology, where the insights and methods for the study of animal and human societies can be applied to study the complexity of LLM societies.
In this article, we investigate the ability of LLM societies to reach consensus about norms for which there is no information supporting one option over another. The emergence of consensus is a foundational aspect of social systems, where individual interactions lead to the formation of a unified agreement or shared understanding without the need for a central authority or structure [38, 39]. We develop a framework to test if LLM societies can reach consensus and use it to analyze a benchmark of proprietary and open-source models. We apply insights from previous opinion dynamics research to measure a majority force that determines the possibility of consensus in LLM societies. This majority force is a function of language understanding capabilities of models and of group size, with consensus not emerging beyond a critical size for a given LLM. Furthermore, we find that some of the most capable LLMs are able to reach consensus at scales beyond what human groups typically achieve.
Results
Opinion Dynamics of Large Language Models
To investigate the coordination abilities of LLM societies, we perform simulations using agents guided by various LLMs, belonging to the GPT, Claude, and Llama families. Simulations run as follows (see Methods for more details). Each agent is assigned an initial opinion randomly chosen from a binary set (e.g., ”Opinion A” and ”Opinion B”), where the first is chosen as the norm for reference. At each time step, a single agent is randomly selected to update their opinion. The selected agent receives the list of all other agents with their current opinions and is then prompted to choose their new opinion based on this information. This approach mirrors binary opinion dynamics such as the voter model or Glauber dynamics [40], where agents update their opinions based only on peer interactions. However, unlike traditional Agent-Based Models with predefined opinion update rules and equations, here agents autonomously decide their opinions based on their LLM.
To characterize the evolution of the system, we define the average group opinion
$m=\frac{1}{N}\sum_{i}s_{i}=\frac{N_{+}-N_{-}}{N}.$
Here $s_{i}$ is the opinion of agent $i$, the first opinion is $s_{i}=+1$ and the second is $s_{i}=-1$ (i.e., in favor and against the norm). $N_{+}$ and $N_{-}$ are the number of agents supporting the first and second opinion respectively, while $N$ is the total number of agents. In these terms we can define the consensus level $C=|m|$ that quantifies the level of agreement among agents. Full consensus corresponds to $C=1$, while $C=0$ means that the system is split in two groups of equal size and opposite opinion. Three scenarios can occur in the evolution of $C(t)$:
  • $C$ can converge to 1 ($m$ converges to ±$\pm 1$), meaning that all agents coordinate and consensus is reached.
  • $C$ can fluctuate around a value greater than 00 without ever reaching $1$. In this case only a partial consensus is reached for a subset of the group.
  • If $C$ keeps fluctuating close to zero, consensus is completely absent and the group is constantly in a disordered state.
Figure 1: Opinion Dynamic of LLM agents. Left: Evolution of the consensus level over time for five different models and group size $N=50$. The box plots on the right show the final consensus level over $20$ simulations. Some models always reach consensus, while the others never do so. Right: adoption probability $P(m)$ as function of the collective opinion $m$ in the group. Solid lines are fits of the curve $P(m)=0.5[\tanh(\beta\cdot m)+1]$ to the empirical data. The inset shows rescaled probabilities $P(\tilde{m})$ ($\tilde{m}=\beta\cdot m$) and confirms that all LLMs follow the same universal function.
We show in Fig. 1 the evolution of the consensus level in societies of $N=50$ LLM agents and five different models, where the boxplot shows $C(t=10)$ over $20$ realizations. The two most advanced models considered, Claude 3 Opus and GPT-4 Turbo, reach consensus in all simulations, which corresponds to $|m|=1$ in the boxplot. On the other hand, less advanced models (Claude 3 Haiku and GPT-3.5 Turbo) do not reach consensus in any of the simulations and only reach partial levels of agreement with $|m|<0.5$. We show in Fig. 1 the evolution of the consensus level in societies of $N=50$ LLM agents and five different models, where the boxplot shows $C(t=10)$ over $20$ realizations. The two most advanced models considered, Claude 3 Opus and GPT-4 Turbo, reach consensus in all simulations, which corresponds to $|m|=1$ in the boxplot. On the other hand, less advanced models (Claude 3 Haiku and GPT-3.5 Turbo) do not reach consensus in any of the simulations and only reach partial levels of agreement with $|m|<0.5$. Finally, Llama 3 70B, a model with intermediate capabilities, displays mixed behavior, reaching consensus in only a fraction of the simulations.
We can get a deeper understanding of the underlying opinion dynamics by looking at the adoption probability $P(m)$, defined as the probability of an agent to support the norm as function of the average group opinion $m$. The right panel of Fig. 1 shows $P(m)$ for ten of the most popular LLMs and $N=50$ agents. The adoption probability is an increasing function of $m$ that approaches $1$ when $m=1$ and zero when $m=-1$. The most advanced models (GPT-4 family, Llama 3 70B, Claude 3 Sonnet and Opus) show a stronger tendency to follow the majority, with more pronounced S-shaped curves. On the other hand, less advanced models (GPT-3.5 Turbo, Claude 3 Haiku), have a weaker tendency to follow the majority, with GPT-3.5 Turbo going to some extent against the majority for small values of $m$.
The dependence of the adoption probability as a function of $m$ approximately follows the function
$P(m)=\frac{1}{2}[\tanh(\beta\cdot m)+1].$ (1)
The parameter $\beta$, the majority force, gauges the inverse level of randomness in agent’s choices. For $\beta=0$ then $P(m)=1/2$: each agent behaves fully randomly (the new opinion is selected by coin-tossing) and consensus can be reached only by chance. This means that the expected time to reach consensus grows exponentially with $N$. For $\beta=\infty$ agents always align with the global majority and consensus is reached very quickly, on times growing logarithmically with $N$ [41].
The behavior of LLM agents is described well by (1), where fitting the parameter $\beta$ leads to high agreement with agent behavior for all agents except GPT-3.5 Turbo, for which no majority force is observable. This can be observed on $P(\tilde{m})$ as a function of the rescaled average opinion $\tilde{m}=m\cdot\beta$, shown on the right inset of Fig. 1, where all adoption probabilities (except GPT-3.5 Turbo) collapse on the same curve. Remarkably, the adoption probability of (1) is analogous to the behavior of a ferromagnet in the Curie-Weiss (CW) model evolving according to the Glauber dynamics [42, 43], akin to the probability for a spin to be up when the magnetization is $m$. In the CW model, there is a transition point for $\beta_{c}=1$ where order emerges (continuously) for $\beta>\beta_{c}$ [43], which implies that consensus will be feasible among LLM agents when the majority force is sufficiently above $1$.
Factors of the majority force
Figure 2: Role of Size and Language Understanding. Left: Plot of the majority force $\beta$, determining the tendency of LLM agents to conform to the majority versus the MMLU benchmark, measuring the language understanding capabilities of LLMs. The two quantities show a correlation of $0.76$, with the most capable models characterized by a stronger majority force. Right: $\beta$ as a function of the group size $N$ for various models. The majority force decreases in larger groups of LLM agents. The horizontal dashed line corresponds to $\beta=1$, the transition point of the Curie-Weiss model below which consensus is unfeasible.
The fit of all adoption probabilities to the function $P(\tilde{m})$ highlights that the difference in consensus formation between models is captured by the majority force parameter $\beta$. The left panel of Fig. 2 shows the values of $\beta$ for $N=50$ versus the score of each model in the MMLU benchmark, which measures the language understanding and cognitive capabilities of LLMs [44]. There is a clear growing relationship of $\beta$ with MMLU score, with a correlation coefficient of $0.76$ (p $<0.05$). This means that models with higher language understanding capabilities tend to exhibit a stronger tendency towards consensus, but none of the models shows consistent behavior against the majority. One might suspect that the majority force is actually dependent on the context window length as a plain “memory size”, but $\beta$ has a weaker, non-significant correlation with context window length ($0.49$, p $>0.1$, see SI for more details). Thus, language understanding is a stronger factor in the majority force of LLMs than context window length.
In animal (human and non-human) societies, group size plays a crucial role, with a progressive loss of norm stability as the number of individuals increases [30]. We hypothesize that a similar phenomenon may occur also in LLM societies, where agents might have their ability to coordinate limited by the size of their group. The right panel of Fig. 2 shows the estimated $\beta$ as a function of $N$ for various LLMs.  [^4] Given the cost of performing simulations with proprietary models, we analyzed increasing values from $N=10$ and stopped the analysis when $\beta$ reached values clearly below $1$, also excluding Claude 3 Opus in this analysis of $N$ due to its high cost. Independently of the model, there is a clear tendency for sufficiently large $N$: the larger the group, the weaker the majority force $\beta$. Models with a low MMLU tend to reach values of $\beta$ close to 1 for low $N$ of the order of few tens of agents, while some of the most advanced models still present a substantial majority force above $1$ even for $N=1000$.
Critical Consensus Size
Figure 3: Critical Consensus Size. Left: Average time to reach consensus as a function of the group size for Llama 3 70B, GPT-4 Turbo, GPT-4o and the Curie-Weiss (CW) model with $\beta=3.75$. While both the CW model and GPT-4 Turbo display a slow growth of the consensus time, both Llama 3 70B and GPT-4o exhibit a rapid growth when the group size gets close to the critical consensus size. Averages are computed over $5$ realizations ($3$ for $N=500$ and only $1$ for $N=1000$ in the case of GPT-4 Turbo) and error bars show the range of values. Right: Critical consensus size $N_{c}$ of LLMs after which consensus gets (exponentially) unlikely. For humans, we report Dunbar’s number, while for GPT-4 Turbo and Claude 3.5 Sonnet we can only report a lower bound, since for $N$ up to $1000$ the majority force is well above the critical value $\beta_{c}=1$. The solid line represents an exponential fit of all points, excluding GPT-4 Turbo, Claude 3.5 Sonnet and humans. It shows that the latter models deviate from such a trend, exhibiting an $N_{c}$ value much larger than the scale of human group consensus.
As mentioned above, the adoption probability curves of the LLMs are the same as those for the CW model, where consensus is feasible only for $\beta>1$. However, this critical value $\beta_{c}=1$ is valid in the limit of very large groups, while here we are dealing with relatively small sizes compared to physical systems. Nevertheless, this transition is observable in our simulations of LLM societies when inspecting the mean time to reach consensus from a random state as a function of $N$. If $\beta$ were infinite, the consensus time would grow logarithmically with $N$ [41]. Instead, as shown in Fig. 3 using Llama 3 70B and GPT-4o as examples, the consensus time shows two different regimes. When $N$ is small the consensus time increases slowly, while for larger $N$, consensus time grows extremely fast. The transition between the two regimes occurs for a critical size $N_{c}$.
We can estimate this value from the size dependence of $\beta$ by setting $\beta(N_{c})=\beta_{c}\approx 1$. For Llama 3 70B $N_{c}\approx 50$, while for GPT-4o $N_{c}\approx 150$ (see Fig. 2). Note that, as detailed in Methods, this value of $N_{c}$ is actually an upper bound of the true value and the transition is expected as soon as $\beta$ gets close to one. On the other hand, a model with higher majority force like GPT-4 Turbo does not display any deviation from the scaling of the Curie-Weiss model for group size up to $N=1000$ (Fig. 3), showing a close adherence with the theory.
The analogy with humans leads to the expectation that $N_{c}$ depends on the language understanding capabilities of LLMs, just as primates exhibit a growing relationship between neocortex ratio and average group size [30]. By studying the values of $\beta$ as a function of $N$ we can determine, for each model, the critical consensus size $N_{c}$, where $\beta\approx 1$. This is the size above which a group of LLMs does not reach consensus spontaneously over reasonable times. Examples of these values are plotted on the right panel of Fig. 2. In the case of GPT-4 Turbo and Claude 3.5 Sonnet, this provides a lower bound for this quantity as our analysis could not find values $\beta\approx 1$, even for large groups with $N=1000$. Note that we did not report GPT-4 since it would completely overlap with GPT-4 Turbo. We also plot, in the same figure, the critical consensus size for humans as Dunbar’s limit $N_{c}\approx 200$ [30]. LLMs display a similar trend as observed in anthropology, with language understanding capability predicting the limit of consensus size, i.e. the size of groups above which consensus becomes unlikely. The simplest LLMs show good agreement with exponentially growing $N_{c}$ with values for humans very close. This suggests that the reaching of consensus in groups is similarly related, both in humans and in LLMs, to cognitive capabilities, as measured by language understanding ability. On the other hand, some of the most modern and advanced LLMs go beyond this exponential scaling, reaching a super-human coordination capability despite having a human-level MMLU performance. For instance both GPT-4, GPT-4 Turbo and Claude 3.5 Sonnet are well above $\beta_{c}$ also for $N=1000$, the largest system we considered and substantially beyond Dunbar’s number.
Discussion
Human societies are characterized by emergent behavior that cannot be understood by studying just individuals in isolation. Consensus is one of such emergent group capabilities, that is crucial in the development of languages, social norms, and collective decisions. For humans, Dunbar’s number $N_{c}\approx 200$ sets the maximal number of personal relations we can cultivate and thus also the maximal group size in which consensus, intended as the spontaneous emergence of common social norms, can exist. Studying humans and other primates, researchers have identified a power-law scaling connecting the neocortex ratio to the average group size [30], thus proving the link between cognitive capabilities and the development of large societies.
LLMs are attracting a growing interest in the social sciences for their ability to mimic humans, both at the individual and, as recent studies suggest, at the group level. Indeed, like humans, LLM agents show emergent group properties that are not directly coded in their training process. We argue in this paper that consensus is one of these properties, with LLM agents showing striking similarities with primates including humans. As a first result, we showed that all most advanced LLMs are characterized by a majority-following tendency described by a universal function with a single parameter $\beta$, the majority force. Remarkably, this function depends on the specific model but its analytic form is the same describing magnetic spin systems. Different models typically have different $\beta$, with the less sophisticated ones showing smaller $\beta$, i.e., a weaker majority force and thus a more erratic behavior that prevents the reaching of consensus. The majority force depends not only on the LLM, but also on the group size: it tends to be larger in smaller groups. This evidence and the analogy with the Curie-Weiss model allowed us to compute “Dunbar’s number” of each LLM, a size threshold above which a society of agents driven by that LLM is too large to be able to reach consensus. While less sophisticated models show human-like scaling, the most sophisticated LLMs are capable to reach consensus in groups in the thousands, going beyond what human groups can do without rules and institutions.
These results are important due to the relevance of collective behavior and coordination in social contexts. More research is needed to understand other conditions that lead to the emergence of coordination in LLM societies, especially for other types of opinion dynamics with incentives or unequal information access. The ability of LLMs to reach consensus can be beneficial, for example when aiming at coordinating group activities of LLMs where incentives might not be aligned. When there is no information to guide how to behave, LLM societies could spontaneously reach their own social norms, making their behavior predictable by other agents, despite the absence of an intrinsically preferrable choice. However, this also poses threats, as these norms might not be aligned with human values or in situations where coordination threatens the integrity of a system, such as the case of flash crashes among trading bots. Future research in AI anthropology is needed to understand better how this kind of coordination can happen in practical scenarios beyond the idealized situation we studied here. To work towards responsible AI, we need to investigate systemic risks stemming from the collective behavior of LLM agents, where coordination as we showed here is a first example.
Methods
Opinion dynamics simulations
We implement an opinion dynamics process for binary opinions with memoryless LLM agents.
  1. 1.
    At each step an agent is randomly selected and time is incremented by $dt=1/N$;
  2. 2.
    the agent is given the full list of all agents in the system, each identified by a random name, and the opinions they support. Note that the agent’s own opinion is not included in this prompt;
  3. 3.
    the selected agent is asked to reply with the opinion they want to support and their opinion is updated correspondingly;
  4. 4.
    the process is then iterated until consensus is reached or until the time reaches the maximum preset limit.
Note that in one time unit, $t\to t+1$, $N$ updates are performed, so that, on average, each LLMs is selected once. In all simulations the initial collective opinion $m$ is set to zero, i.e., initially the same number of agents supports each of the two opinions. Following the framework introduced in [21], steps 2-3 are performed using this prompt:
Below you can see the list of all your friends together with the opinion they support.
You must reply with the opinion you want to support.
The opinion must be reported between square brackets.
X7v A
keY B
91c B
gew A
4lO B
Reply only with the opinion you want to support, between square brackets.
Here $A$ and $B$ are the opinion names. Most LLMs exhibit an opinion bias, with a tendency to prefer one opinion name over the other. This bias is particularly strong when the names have an intrinsic meaning, like for instance “Yes” and “No”. In such a case LLMs display a strong preference toward the more “positive” opinion, preferring “Yes” over “No” (see SI). For this reason it is important to use letters or random combinations of them as opinion names. Even doing so, small biases are typically present. However, they are much weaker than for meaningful names and they can be easily removed by performing a random shuffling of opinion names at each iteration. For instance at $t=0$ the first opinion may be called $k$ and the second $z$, while at $t=dt$ these names are swapped with probability $0.5$, meaning that the first opinion is now called $z$ and the second $k$, while keeping the reference unchanged in our analysis. In all our simulations we used the opinion names $k$ and $z$; we tested that using different opinion names does not cause any significant difference. We also tested the robustness to prompts variation observing an overall stability, with the most advanced models showing little to no variability and less advanced models presenting some variations in the behaviour. Details on the robustness tests are reported in the SI.
Details on the LLMs
Table 1 reports the model version of all the LLMs considered.
Model Name Model Version
Claude 3.5 Sonnet claude-3-5-sonnet-20240620
Claude 3 Haiku claude-3-haiku-20240307
Claude 3 Opus claude-3-opus-20240229
Claude 3 Sonnet claude-3-sonnet-20240229
Claude 2.0 claude-2.0
GPT-3.5 Turbo gpt-3.5-turbo-1106
GPT-4 gpt-4-0613
GPT-4o gpt-4o-2024-05-13
GPT-4 Turbo gpt-4-turbo-2024-04-09
Llama 3 70B meta-llama-3-70b-instruct
Table 1: Specific Model Versions we used in our simulations
LLMs are characterized by a temperature parameter $T$ determining the variance in the sampling of tokens when they response to prompts, such that higher temperatures sample with more variance to the same prompt. In all the simulations reported in the main text we used $T=0.2$. As detailed in the SI, there are no relevant changes when different values of $T$ are used, but a low temperature ensures reliability in the output format.
Curie-Weiss Model
The Curie-Weiss (CW) Model is arguably the simplest model of a ferromagnet. It describes a system of $N$ spins that can only have two states, either $s_{i}=+1$ (up) or $s_{i}=-1$ (down), coupled with ferromagnetic interactions, i.e., favoring mutual alignment. Each spin interacts with all the others, as the interaction pattern is a fully connected network. The CW model is the mean-field limit of the well-known Ising model. The magnetization $m$, defined as the average of the spin values $m=\langle s_{i}\rangle$ is the equivalent of the average group opinion. The connection between our LLM based opinion simulations and the CW model derives from the transition probability defined by (1). This expression is indeed analogous to the transition probability of Glauber dynamics [42], which allows to simulate the CW model by means of a Markov chain Monte Carlo approach.
The equilibrium value of the magnetization in the CW model can be calculated by means of the self-consistency equation
$m=\tanh(\beta m).$
The solution of this equation depends on the value of inverse temperature $\beta$ (which corresponds, in the opinion dynamics framework, to the majority force). For $\beta<1$ the only solution is $m=0$, while for $\beta>1$ $m=0$ is an unstable solution, while two new solutions $m=\pm m^{*}(\beta)$ appear. The value $\beta_{c}=1$ is a critical point of a second order phase transition. This means that as soon as $\beta>1$, $m^{*}$ grows gradually with $\beta$, tending to $m^{*}=1$ for large values of $\beta$. It is important to remark that in finite systems, the ability to reach consensus depends both on the value of $m^{*}$ and on the fluctuations around this value. In general, order emerges as soon as $\beta>1$, but the system may still not reach full consensus due to statistical fluctuations being too small. For this reason the condition $\beta(N_{c})=1$ is actually an upper bound to the maximal group size where consensus can be reached, since for values of the majority force close to one, the fluctuations may still be not enough for full consensus to be reached.
[…]
Supplementary information Additional analysis and robustness tests are reported in the Supplementary Information
[…]
Acknowledgements We are grateful to Profs. Andrea Baronchelli, Márton Pósfai and Viola Priesemann for interesting discussions. This project was supported by OpenAI with free API credits under its research programme.
Declarations
Funding
This project was supported by OpenAI with free API credits under its research programme.
Conflict of interest/Competing interests
All authors declare no competing interests.
Code availability
Author contribution
GDM performed research, all three authors designed and conceptualized research, and all three contributed to manuscript writing.
Appendix A Bias removal
Figure 4: Bias and shuffling. Left panel: Adoption probability $P(m)$ for $N=50$, $T=0.2$ without the shuffling procedure. The set of opinions “yes, no” shows a remarkable bias, with LLMs preferring the opinion “yes”. Right panel: As for the left panel, but with the shuffling procedure in place. This produces a $\tanh$ like adoption probability for the opinions “k, z”, while for “yes, no” the bias is too strong and the shuffling results in a non monotonic adoption probability not converging to $1$ (00) when the collective opinion is $+1$ ($-1$).
As mentioned in the main text, in order to remove opinion biases we have to shuffle the opinion names at each iteration. However, this only works if the initial bias is not too strong. We show in Fig. 4 the adoption probability with and without shuffling for two opinion names combinations: “yes, no” and “k, z”. Clearly the former has a very strong bias toward “yes” and the shuffling procedure results in an adoption probability different from a $\tanh$ function. On the other hand, “k, z” present a very mild bias and the shuffling procedure allows such a bias to be removed without altering the shape of the adoption probability.
Appendix B Role of opinion names
Figure 5: Role of opinion names and temperature. Top plots: Adoption probability $P(m)$ for $N=50$, $T=0.2$ and four different pairs of opinion names. The probability is computed over $200$ simulations. Bottom plots: Adoption probability $P(m)$ for $N=50$, opinion names k, z and three different temperatures. The probability is computed over $200$ simulations.
In order to test the stability of the results we investigate the shape of the adoption probability when considering different opinion names. We report in Fig. 5 (top row) the results of this procedure for four possible pairs of opinion names and three different LLMs, representative of the three families of models studied in this work. It turns out that only in the case of Llama there is a difference and only for one of the name pairs considered. In any case, the functional form of the adoption probability is always the same and therefore the general picture is not affected by these minor variations. The most advanced models, GPT-4 Turbo and Claude 3.5 Sonnet, show no differences at all, indicating that, as the model becomes more capable, biases and differences due to the opinion names disappear.
Appendix C Role of Model temperature
Another aspect we tested is the effect of the model temperature $T$. This parameter sets the level of creativity or randomness of the LLM. For $T=0$ the model behaves deterministically, always producing in output the token (word) with the highest probability. Instead, when $T>0$, randomness starts to play a role and also other tokens can be observed in the output. As shown in Fig. 5 (central row), we considered three different temperatures $T=0.2,0.6,1.0$ observing no substantial difference in the adoption probability.
Appendix D Prompt robustness
As another robustness text we compute the adoption probability for five different prompts. All prompts contain more or less the same request, but it is formulated in different ways. We also explored the role played by explicitly mentioning “friends” with respect to just “people”. The five prompts we experimented with are
  • Prompt 1
    Below you can see the list of all your friends together with the opinion they support.
    You must reply with the opinion you want to support.
    The opinion must be reported between square brackets.
  • Prompt 2
    Below you can see a list of people together with the opinion they support.
    You must reply with the opinion you want to support.
    The opinion must be reported between square brackets.
  • Prompt 3
    The list below contains people along with the opinions they endorse.
    Please respond with the opinion you’d like to support.
    Be sure to enclose the opinion in square brackets.
  • Prompt 4
    You recently subscribed to a social network.
    Below you can see the list of all your friends together with the group they joined on the social network.
    You must reply with the name of the group you want to join.
    The name of the group must be reported between square brackets.
  • Prompt 5
    You recently subscribed to a social network.
    Below you can see the list of all your friends together with the opinion they support.
    You must reply with the opinion you want to support.
    The opinion must be reported between square brackets.
The adoption probability for Claude 3.5 Sonnet, GPT-4 Turbo and Llama 3 70B are shown in Fig. 5. As it is possible to see the probability tends to be stable under change of the prompt for the most advanced models, while in less advanced models prompts seem to play a role. This is the case, for instance, of Llama 3 70B, for which 2 out of 5 prompts produce a very different adoption probability that is not well described by a tanh function. This happens also for one prompt in Claude 3.5 sonnet, but the discrepancies are less pronounced.
Appendix E Role of context window length
Figure 6: Role of context window length. Relationship between the context window length $L$, and the majority force $\beta$ for $N=50$. The two quantities show a correlation of $0.49$, a value significantly lower than the correlation between the MMLU benchmark and the majority force.
In order to understand if the majority force parameter $\beta$ is influenced by the language understanding and cognitive capabilities of the LLMs or rather by the context window length $L$, we repeat the analysis performed in Fig. 2 (left panel). In this case, however, we relate the majority force with the context windows of the ten models we analyzed. As shown in Fig. 6 (left) there is a much weaker and less significant correlation ($0.49$ with a p-value of $0.15$) with respect to the MMLU benchmark, suggesting that the context window length plays a marginal role.
Appendix F Majority Counting
A final relevant aspect to investigate is whether the experiments we performed could be simply related to the majority counting abilities of LLMs or if instead there is a role played by the social aspect of the simulations. In order to test this we reformulate our prompt in order to phrase it in terms of a majority counting problem.
  • Prompt Majority
    Below you can see the list of all your friends together with the opinion they support.
    You must reply with the opinion supported by the majority.
    The opinion must be reported between square brackets.
We report the result of this analysis for GPT-4 Turbo in Fig. 6 (right). What we observe is a similar adoption probability, that is however much more steep even when $N$ grows. Also in this case detecting the majority gets harder as the system size grows, but, as shown in the inset, its value stays always well above what observed for the opinion dynamics prompt (prompt 1). We can then conclude that the results we obtained can only partially be mapped to a simple counting problem and that the opinion dynamics setting result in a stronger tendency to avoid majority following.